OpenStudy (anonymous):

Which graph represents a phase shift of π/2 units left for the graph of y = sinx ?

OpenStudy (anonymous):

A. y = sin (x - (π/2)) B. y = cos x C. y = sec x D. y = csc x

terenzreignz (terenzreignz):

Well, only one of them looks so unlike the others that it most probably has the phase shift :P

OpenStudy (anonymous):

But its not A.

terenzreignz (terenzreignz):

But in all seriousness, the phase-shift is that constant added to the x (inside the trig function) \[\Large y = \sin(px+q)\]In that case, there is a phase shift of \(\huge \frac{q}p\) units to the left...

terenzreignz (terenzreignz):

whoops... sorry about that...

OpenStudy (anonymous):

Its okay! The only thing I can think of is co"sine"

terenzreignz (terenzreignz):

I guess you use this fact... \[\Large \sin\left(\frac\pi 2 - x\right)=\cos(x)\]

OpenStudy (anonymous):

Yup, B.

terenzreignz (terenzreignz):

I'm actually now confused... oh well, it usually comes with weekends... LOL Signing off now... ------------------------- Terence out